OpenStudy (anonymous):

Choose 2 axioms that allows 22 + (m + 8) to be written as m + 30 commutative - addition distributive associative - multiplication symmetric commutative - multiplication associative - addition identity - addition

OpenStudy (anonymous):

commutative for addition \[22+(m+8)=(m+8)+22\] then associative \[(m+8)+22=m+(8+22)\] \[=m+30\]

OpenStudy (anonymous):

don't even need that, just do the associative property of addition and you're done. \[22 + (m + 8) = m + (22+8)\]

OpenStudy (anonymous):

unless you need to choose two of them that could do it, then satellite's right.

OpenStudy (anonymous):

not to get picky but without the commutative law you cannot get from \[22+(m+8)\] to \[m+(22+8)\]